10. What is the risk premium for Stock C, given it has a beta of 2, and the market risk premium of 12%.8. What is the expected return for a stock that always pays a $3 annual dividend and sells now for $91?
Answer #1

10. Risk premium for stock C =beta*market risk =2*12% =24%

8. Expected return =dividend/price =3/91 =3.30%